# A posteriori error estimation for an interior penalty virtual element method of Kirchhoff plates

### v2

@@ 4 Enriching operator
$h_{z}|\nabla(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)|$ $=h_{z}|\nabla\Pi_{h}^{\Delta}v(z)|\leq h_{K}\|\nabla\Pi_{h}^{\Delta}v\|_{\infty,e}\lesssim h_{e}^{1/2}\|\nabla\Pi_{h}^{\Delta}v\|_{0,e}\lesssim h_{e}^{-1/2}\|\Pi_{h}^{\Delta}v\|_{0,e}$
$=h_{e}^{-1/2}\|\Pi_{h}^{\Delta}v-v\|_{0,e}\lesssim|v-\Pi_{h}^{\Delta}v|_{1,K}+h_{K}^{-1}\|v-\Pi_{h}^{\Delta}v\|_{0,K}.$
Therefore, for any vertex $z$ of $\mathcal{T}_{h}$, by using the norm equivalence (3.17), we obtain
$\Big{(}h_{K}^{-1}h_{z}|\nabla(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)\Big{)}^{2}\lesssim S^{K}(v-\Pi_{h}^{\Delta}v,v-\Pi_{h}^{\Delta}v)+\sum_{e_{i}}\Big{(}\frac{1}{|e_{i}|^{3}}\|[v]\|_{0,e_{i}}^{2}+\frac{1}{|e_{i}|}\|[\partial_{\bm{n}_{e_{i}}}v]\|_{0,e_{i}}^{2}\Big{)}.$
This completes the proof. ∎
@@
Our result aligns with the finite element methods since the first term on the right-hand side of (4.2) vanishes in this case (see [9, 8, 12] for instance).
For the medius error analysis, we should bound the (semi-)norm of $v-E_{h}v$ by $|v|_{2,h}$. Here, we present the result for $\|v-E_{h}v\|_{0,K}$ as an example, even though this estimate will not be needed in the subsequent analysis.
Lemma 4.2.
For any $v\in V_{h}$ and $E_{h}v$ defined above, there holds
$\sum\limits_{K\in\mathcal{T}_{h}}h_{K}^{-4}\|v-E_{h}v\|_{0,K}^{2}\lesssim|v|_{2,h}^{2}.$
Proof.
From the triangle inequality
$\|v-E_{h}v\|_{0,K}\leq\|v-\Pi_{h}^{\Delta}v\|_{0,K}+\|\Pi_{h}^{\Delta}v-E_{h}v\|_{0,K}$
and the standard estimate
$h_{K}^{-2}\|v-\Pi_{h}^{\Delta}v\|_{0,K}\lesssim|v|_{2,K},$
it remains to bound the term $\|\Pi_{h}^{\Delta}v-E_{h}v\|_{0,K}$. By the norm equivalence,
$h_{K}^{-2}\|\Pi_{h}^{\Delta}v-E_{h}v\|_{0,K}\simeq h_{K}^{-1}\|\bm{\chi}^{c}(\Pi_{h}^{\Delta}v-E_{h}v)\|_{l^{2}}.$
Let $z$ be an interior vertex. As in the proof of Lemma 4.1, one only needs to consider $|(\Pi_{h}^{\Delta}v)_{1}(z)-(\Pi_{h}^{\Delta}v)_{2}(z)|$. Noting that $v$ is continuous at $z$ (i.e., $v_{1}(z)=v_{2}(z)$), by applying the Sobolev inequality (2.6), we obtain
$|(\Pi_{h}^{\Delta}v)_{1}(z)-(\Pi_{h}^{\Delta}v)_{2}(z)|$ $\leq|(\Pi_{h}^{\Delta}v)_{1}(z)-v_{1}(z)|+|v_{2}(z)-(\Pi_{h}^{\Delta}v)_{2}(z)|$
$\lesssim\sum_{K=K_{1},K_{2}}\Big{(}h_{K}|v-\Pi_{h}^{\Delta}v|_{2,K}+|v-\Pi_{h}^{\Delta}v|_{1,K}+h_{K}^{-1}\|v-\Pi_{h}^{\Delta}v\|_{0,K}\Big{)}.$
The proof is finished by using the error estimate for $\Pi_{h}^{\Delta}$. ∎
@@ 5 A posteriori error estimation
$\eta_{0,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}}\eta_{0,e}^{2}\Big{)}^{1/2},\qquad\eta_{0,e}=\frac{\lambda_{e}^{3/2}}{|e|^{3/2}}\|[u_{h}]\|_{0,e},$
$\eta_{1,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}}\eta_{1,e}^{2}\Big{)}^{1/2},\qquad\eta_{1,e}=\frac{\lambda_{e}^{1/2}}{|e|^{1/2}}\Big{\|}\Big{[}\frac{\partial u_{h}}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e},$
$\eta_{2,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}^{0}}\eta_{2,e}^{2}\Big{)}^{1/2},\qquad\eta_{2,e}=|e|^{1/2}\Big{\|}\Big{[}\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}_{e}^{2}}\Big{]}\Big{\|}_{0,e},$
$\eta_{3,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}^{0}}\eta_{3,e}^{2}\Big{)}^{1/2},\qquad\eta_{3,e}=|e|^{3/2}\Big{\|}\Big{[}\frac{\partial(\Delta\Pi_{h}^{\Delta}u_{h})}{\partial\bm{n}_{e}}+\frac{\partial^{3}(\Pi_{h}^{\Delta}u_{h})}{\partial\bm{n}_{e}\partial\bm{t}_{e}^{2}}\Big{]}\Big{\|}_{0,e},$
$\eta_{4,K}=\big{(}S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},u_{h}-\Pi_{h}^{\Delta}u_{h})\big{)}^{1/2},$
@@
$\eta_{K}(u_{h})=\Big{(}\sum\limits_{i=0}^{6}\eta_{i,K}^{2}\Big{)}^{1/2}$
and
$\eta(u_{h})=\eta(u_{h},\mathcal{T}_{h})=:\Big{(}\sum\limits_{K\in\mathcal{T}_{h}}\eta_{K}^{2}(u_{h})\Big{)}^{1/2}.$
@@ 5.1   Upper bound
$\lesssim|u-E_{h}u_{h}|_{2}+\Big{(}\sum_{K\in\mathcal{T}_{h}}S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},u_{h}-\Pi_{h}^{\Delta}u_{h})$
$\qquad+\sum_{e\in\mathcal{E}_{h}}\frac{1}{|e|^{3}}\|[u_{h}]\|_{0,e}^{2}+\sum\limits_{e\in\mathcal{E}_{h}}\frac{1}{h_{e}}\Big{\|}\Big{[}\frac{\partial u_{h}}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\Big{)}^{1/2}+J_{1}(e_{h},e_{h})^{1/2}$
$\lesssim|u-E_{h}u_{h}|_{2}+\Big{(}\sum_{K\in\mathcal{T}_{h}}(\eta_{0,K}^{2}+\eta_{1,K}^{2}+\eta_{4,K}^{2})\Big{)}^{1/2}.$
@@
Let $\phi=u-E_{h}u_{h}$ and denote by $I_{h}:H^{2}(\Omega)\to V_{h}$ the interpolation operator. According to the definitions of the continuous and discrete variational problems, one gets
@@
For the term $I_{1}$, using the Cauchy-Schwarz inequality and the estimate (4.2), one gets
@@
$\lesssim\Big{(}\sum_{e\in\mathcal{E}_{h}}\eta_{1,e}^{2}\Big{)}^{1/2}|\phi|_{2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}(\eta_{0,K}^{2}+\eta_{1,K}^{2}+\eta_{4,K}^{2})\Big{)}^{1/2}|\phi|_{2}.$
For the term $I_{2}$, using the integration by parts we can split it into four terms as
$I_{2}$ $=\sum_{K\in\mathcal{T}_{h}}(f_{h}-\Delta^{2}\Pi_{h}^{\Delta}u_{h},\phi-I_{h}\phi)_{K}$
$\qquad+\sum_{K\in\mathcal{T}_{h}}\Big{(}\frac{\partial(\Delta\Pi_{h}^{\Delta}u_{h})}{\partial\bm{n}}+\frac{\partial^{3}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}\partial\bm{t}^{2}},\phi-I_{h}\phi\Big{)}_{\partial K}$
$\qquad-\sum_{K\in\mathcal{T}_{h}}\Big{(}\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}^{2}},\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\Big{)}_{\partial K}$
$\qquad+\sum_{K\in\mathcal{T}_{h}}\sum\limits_{z\in\mathcal{V}(K)}\Big{[}\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{t}\partial\bm{n}}\Big{]}(z)(\phi-I_{h}\phi)(z)$
$=:I_{2,1}+I_{2,2}+I_{2,3}+I_{2,4}.$
For $I_{2,1}$, by the interpolation error estimate,
$I_{2,1}$ $=\sum_{K\in\mathcal{T}_{h}}(f_{h}-\Delta^{2}\Pi_{h}^{\Delta}u_{h},\phi-I_{h}\phi)_{K}\leq\sum_{K\in\mathcal{T}_{h}}\|f_{h}-\Delta^{2}\Pi_{h}^{\Delta}u_{h}\|_{0,K}\|\phi-I_{h}\phi\|_{0,K}$
$\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{2}\|f_{h}-\Delta^{2}\Pi_{h}^{\Delta}u_{h}\|_{0,K}|\phi|_{2,K}\leq\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{6,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$
For the sake of brevity, we set $Q_{3}(u_{h})=\frac{\partial(\Delta\Pi_{h}^{\Delta}u_{h})}{\partial\bm{n}}+\frac{\partial^{3}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}\partial\bm{t}^{2}}$ in $I_{2,2}$. From the equality (2.2), one gets
@@
$I_{2,4}=\sum_{K\in\mathcal{T}_{h}}\sum\limits_{z\in\mathcal{V}(K)}\Big{[}\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{t}\partial\bm{n}}\Big{]}(z)(\phi-I_{h}\phi)(z)=0.$
Hence, it follows from the Cauchy-Schwarz inequality, the trace inequality and the interpolation error estimates that
@@
For $I_{2,3}$, noting that $[\frac{\partial\phi}{\partial\bm{n}}]_{e}=0$, from (2.2), we have
$I_{2,3}$ $=-\sum_{K\in\mathcal{T}_{h}}\Big{(}\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}^{2}},\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\Big{)}_{\partial K}$
$=-\sum_{e\in\mathcal{E}_{h}^{0}}\Big{(}([\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}^{2}}],\{\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\})_{e}-\sum_{e\in\mathcal{E}_{h}}(\{\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}^{2}}\},[\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}])_{e}\Big{)}$
$=-\sum_{e\in\mathcal{E}_{h}^{0}}([\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}^{2}}],\{\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\})_{e}-J_{2}(u_{h},I_{h}\phi).$
Similarly, by using the Cauchy-Schwarz inequality, the trace inequality and the interpolation error estimates, we can get
$I_{2,3}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{2,K}^{2}\Big{)}^{1/2}|\phi|_{2}-J_{2}(u_{h},I_{h}\phi).$
@@
$I_{2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}(\eta_{2,K}^{2}+\eta_{3,K}^{2}+\eta_{6,K}^{2})\Big{)}^{1/2}|\phi|_{2}-J_{2}(u_{h},I_{h}\phi).$
For the term $I_{3}$, by the norm equivalence, we obtain
@@
$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}^{2}\Big{)}^{1/2}\Big{(}\sum_{K\in\mathcal{T}_{h}}\big{(}|\phi|_{2,K}^{2}+|\phi-{I_{h}}\phi|_{2,K}^{2}\big{)}\Big{)}^{1/2}$
$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},\,u_{h}-\Pi_{h}^{\Delta}u_{h})\Big{)}^{1/2}|\phi|_{2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{4,K}^{2}(u_{h})\Big{)}^{1/2}|\phi|_{2}.$
@@
For $J_{1}(u_{h},I_{h}\phi)$, by the interpolation error estimate, we have
$J_{1}(u_{h},I_{h}\phi)\lesssim(\sum_{e\in\mathcal{E}_{h}}\eta_{1,e}^{2})^{1/2}\Big{(}\sum_{e\in\mathcal{E}_{h}}|e|^{-1}\Big{\|}\Big{[}\frac{\partial(I_{h}\phi-\phi)}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\Big{)}^{1/2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{1,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$
For $J_{3}(u_{h},I_{h}\phi)$, the boundedness of $\Pi_{h}^{\Delta}$ gives
$J_{3}(u_{h},I_{h}\phi)$ $=\Big{|}\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\Delta}I_{h}\phi}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial u_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s\Big{|}\lesssim(\sum_{e\in\mathcal{E}_{h}}\eta_{2,e}^{2})^{1/2}\Big{(}\sum_{K\in\mathcal{T}_{h}}|\Pi_{h}^{\Delta}\phi|_{2,K}^{2}\Big{)}^{1/2}$
$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{2,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$ (5.5)

### v4

@@ 4 Enriching operator
$h_{z}|\nabla(\Pi_{h}^{\nabla}v-E_{h}v)|_{K}(z)|$ $=h_{z}|\nabla\Pi_{h}^{\nabla}v(z)|\leq h_{K}\|\nabla\Pi_{h}^{\nabla}v\|_{\infty,e}\lesssim h_{e}^{-1/2}\|\Pi_{h}^{\nabla}v\|_{0,e}$
$=h_{e}^{-1/2}\|\Pi_{h}^{\nabla}v-v\|_{0,e}\lesssim|v-\Pi_{h}^{\nabla}v|_{1,K}+h_{K}^{-1}\|v-\Pi_{h}^{\nabla}v\|_{0,K}.$
Therefore, for any vertex $z$ of $\mathcal{T}_{h}$, by using norm equivalence (3.20) and (3.21), we have
$\Big{(}h_{K}^{-1}h_{z}|\nabla(\Pi_{h}^{\nabla}v-E_{h}v)|_{K}(z)\Big{)}^{2}\lesssim\sum_{e_{i}}\sum_{K\in\mathcal{T}_{e_{i}}}h_{K}^{-2}S^{K}(v-\Pi_{h}^{\nabla}v,v-\Pi_{h}^{\nabla}v)+\sum_{e_{i}}\frac{1}{|e_{i}|}\|[\partial_{\bm{n}_{e_{i}}}v]\|_{0,e_{i}}^{2}.$
The above arguments yield the first estimate (4.2).
Step 2: Let $e$ be an edge shared by two elements $K^{-}$ and $K^{+}$. By the trace inequality (2.5) and the inverse inequality (3.6), one gets
$\|[v]\|_{0,e}=\|[v-I_{h}^{c}v]\|_{0,e}\lesssim\sum_{K=K^{-},K^{+}}h_{K}^{-1/2}\|v-I_{h}^{c}v\|_{0,K}.$
Noting that $\bm{\chi}(v)=\bm{\chi}(I_{h}^{c}v)$, we apply the norm equivalence (3.21) and the norm equivalence for $H^{1}$-conforming virtual elements (cf. [21]) to get
$\|v-I_{h}^{c}v\|_{0,K}\leq\|v-\Pi_{h}^{\nabla}v\|_{0,K}+\|\Pi_{h}^{\nabla}v-I_{h}^{c}v\|_{0,K}\lesssim h_{K}\|\bm{\chi}(v-\Pi_{h}^{\nabla}v)\|_{\ell^{2}},$
which yields the second estimate (4.3).
Step 3: The third estimate follows by considering the triangle inequality
$\Big{\|}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\Big{\|}\leq\Big{\|}\Big{[}\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big{]}\Big{\|}+\Big{\|}\Big{[}\frac{\partial(v-\Pi_{h}^{\nabla}v)}{\partial\bm{n}_{e}}\Big{]}\Big{\|},$ (4.7)
with the details omitted. ∎
@@
Our result aligns with the traditional finite element methods on a triangular mesh since the first term on the right-hand side of (4.2) vanishes when $v$ is replaced by a polynomial of degree $\leq k$ and the remaining parts are the jumps accounting for the discontinuities (see [9, 8, 12] for instance). We remark that the second estimate (4.3) implies that the interior penalty virtual element can be treated as a continuous element in the a posteriori estimate. This property is referred to as the quasi-$C^{0}$ continuity in this article.
@@ 5 A posteriori error estimation
$\eta_{1,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}}\eta_{1,e}^{2}\Big{)}^{1/2},\qquad\eta_{1,e}=\frac{1}{|e|^{1/2}}\Big{\|}\Big{[}\frac{\partial\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e},$
$\eta_{2,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}^{0}}\eta_{2,e}^{2}\Big{)}^{1/2},\qquad\eta_{2,e}=|e|^{1/2}\Big{\|}\Big{[}\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}_{e}^{2}}\Big{]}\Big{\|}_{0,e},$
$\eta_{3,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}^{0}}\eta_{3,e}^{2}\Big{)}^{1/2},\qquad\eta_{3,e}=|e|^{3/2}\Big{\|}\Big{[}\frac{\partial(\Delta\Pi_{h}^{\nabla}u_{h})}{\partial\bm{n}_{e}}+\frac{\partial^{3}(\Pi_{h}^{\nabla}u_{h})}{\partial\bm{n}_{e}\partial\bm{t}_{e}^{2}}\Big{]}\Big{\|}_{0,e},$
$\eta_{4,K}=h_{K}^{-1}\big{(}S^{K}(u_{h}-\Pi_{h}^{\nabla}u_{h},u_{h}-\Pi_{h}^{\nabla}u_{h})\big{)}^{1/2},$
@@
$\eta_{K}(u_{h})=\Big{(}\sum\limits_{i=1}^{6}\eta_{i,K}^{2}\Big{)}^{1/2},\quad\eta(u_{h})=\eta(u_{h},\mathcal{T}_{h})=\Big{(}\sum\limits_{K\in\mathcal{T}_{h}}\eta_{K}^{2}(u_{h})\Big{)}^{1/2}.$
Note that we do not include $\eta_{0,K}$ in the error estimator, where $\eta_{0,K}$ is defined in (3.15).
@@ 5.1   Upper bound
$\lesssim|u-E_{h}u_{h}|_{2}+\Big{(}\sum_{K\in\mathcal{T}_{h}}S^{K}(u_{h}-\Pi_{h}^{\nabla}u_{h},u_{h}-\Pi_{h}^{\nabla}u_{h})$
$\qquad+\sum\limits_{e\in\mathcal{E}_{h}}\frac{1}{|e|}\Big{\|}\Big{[}\frac{\partial\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\Big{)}^{1/2}+J_{1}(e_{h},e_{h})^{1/2}$
$\lesssim|u-E_{h}u_{h}|_{2}+\Big{(}\sum_{K\in\mathcal{T}_{h}}(\eta_{1,K}^{2}+\eta_{4,K}^{2})\Big{)}^{1/2}.$
@@
Let $\phi=u-E_{h}u_{h}$. According to the definitions of the continuous and discrete variational problems, one gets
@@
For the term $I_{1}$, using the Cauchy-Schwarz inequality and the estimate in Lemma 4.1, one gets
@@
$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}(\eta_{1,K}^{2}+\eta_{4,K}^{2})\Big{)}^{1/2}|\phi|_{2}.$
For the term $I_{2}$, using the integration by parts we can split it into several terms as
$I_{2}$ $=\sum_{K\in\mathcal{T}_{h}}a^{K}(\Pi_{h}^{\nabla}u_{h}-\Pi_{h}^{\Delta}u_{h},\phi-I_{h}\phi)+(f_{h},\phi-I_{h}\phi)-\sum_{K\in\mathcal{T}_{h}}a^{K}(\Pi_{h}^{\nabla}u_{h},\phi-I_{h}\phi)$
$=\sum_{K\in\mathcal{T}_{h}}a^{K}(\Pi_{h}^{\nabla}u_{h}-\Pi_{h}^{\Delta}u_{h},\phi-I_{h}\phi)$
$\qquad+\sum_{K\in\mathcal{T}_{h}}(f_{h}-\Delta^{2}\Pi_{h}^{\nabla}u_{h},\phi-I_{h}\phi)_{K}$
$\qquad+\sum_{K\in\mathcal{T}_{h}}\Big{(}\frac{\partial(\Delta\Pi_{h}^{\nabla}u_{h})}{\partial\bm{n}}+\frac{\partial^{3}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}\partial\bm{t}^{2}},\phi-I_{h}\phi\Big{)}_{\partial K}$
$\qquad-\sum_{K\in\mathcal{T}_{h}}\Big{(}\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}^{2}},\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\Big{)}_{\partial K}$
$\qquad+\sum_{K\in\mathcal{T}_{h}}\sum\limits_{z\in\mathcal{V}(K)}\Big{[}\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{t}\partial\bm{n}}\Big{]}(z)(\phi-I_{h}\phi)(z)$
$=:I_{2,0}+I_{2,1}+I_{2,2}+I_{2,3}+I_{2,4}.$
For $I_{2,0}$, we obtain from the boundedness of $\Pi_{h}^{\Delta}$ in Lemma 3.3, the interpolation error estimate (3.8), the inverse inequality (3.6) and the norm equivalence (3.20) that
$I_{2,0}$ $=\sum_{K\in\mathcal{T}_{h}}a^{K}(\Pi_{h}^{\nabla}u_{h}-\Pi_{h}^{\Delta}u_{h},\phi-I_{h}\phi)\leq\sum_{K\in\mathcal{T}_{h}}|\Pi_{h}^{\nabla}u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}|\phi-I_{h}\phi|_{2,K}$
$\lesssim\sum_{K\in\mathcal{T}_{h}}|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{2,K}|\phi|_{2,K}\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{-1}|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{1,K}|\phi|_{2,K}$
$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{4,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$
For $I_{2,1}$, by the interpolation error estimate (3.8),
$I_{2,1}$ $=\sum_{K\in\mathcal{T}_{h}}(f_{h}-\Delta^{2}\Pi_{h}^{\nabla}u_{h},\phi-I_{h}\phi)_{K}\leq\sum_{K\in\mathcal{T}_{h}}\|f_{h}-\Delta^{2}\Pi_{h}^{\nabla}u_{h}\|_{0,K}\|\phi-I_{h}\phi\|_{0,K}$
$\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{2}\|f_{h}-\Delta^{2}\Pi_{h}^{\nabla}u_{h}\|_{0,K}|\phi|_{2,K}\leq\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{6,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$
For the sake of brevity, we set $Q_{3}(u_{h})=\frac{\partial(\Delta\Pi_{h}^{\nabla}u_{h})}{\partial\bm{n}}+\frac{\partial^{3}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}\partial\bm{t}^{2}}$ in $I_{2,2}$. From Eq. (2.2) one gets
@@
$I_{2,4}=\sum_{K\in\mathcal{T}_{h}}\sum\limits_{z\in\mathcal{V}(K)}\Big{[}\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{t}\partial\bm{n}}\Big{]}(z)(\phi-I_{h}\phi)(z)=0.$
Hence, it follows from the Cauchy-Schwarz inequality, the trace inequality (2.5) and the interpolation error estimates (3.8) that
@@
For $I_{2,3}$, noting that $[\frac{\partial\phi}{\partial\bm{n}}]_{e}=0$, from (2.2) and the definition of $V_{k}(K)$, we have
$I_{2,3}$ $=-\sum_{K\in\mathcal{T}_{h}}\Big{(}\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}^{2}},\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\Big{)}_{\partial K}$
$=-\sum_{e\in\mathcal{E}_{h}^{0}}\Big{(}([\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}^{2}}],\{\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\})_{e}-\sum_{e\in\mathcal{E}_{h}}(\{\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}^{2}}\},[\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}])_{e}\Big{)}$
$=-\sum_{e\in\mathcal{E}_{h}^{0}}([\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}^{2}}],\{\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\})_{e}-J_{2}(u_{h},I_{h}\phi),$ (5.5)
which gives
$I_{2,3}+J_{2}(u_{h},I_{h}\phi)=-\sum_{e\in\mathcal{E}_{h}^{0}}([\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}^{2}}],\{\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\})_{e}.$
Similarly, by using the Cauchy-Schwarz inequality, the trace inequality (2.5) and the interpolation error estimates (3.8), we can get
$I_{2,3}+J_{2}(u_{h},I_{h}\phi)\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{2,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$
@@
$I_{2}+J_{2}(u_{h},I_{h}\phi)\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}(\eta_{2,K}^{2}+\eta_{3,K}^{2}+\eta_{4,K}^{2}+\eta_{6,K}^{2})\Big{)}^{1/2}|\phi|_{2}.$
For the term $I_{3}$, by the norm equivalence (3.19) and (3.20), the boundedness of $\Pi_{h}^{\Delta}$ in Lemma 3.3, the inverse inequality (3.6), we obtain
@@
$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}^{2}\Big{)}^{1/2}\Big{(}\sum_{K\in\mathcal{T}_{h}}\big{(}|\phi|_{2,K}^{2}+|\phi-{I_{h}}\phi|_{2,K}^{2}\big{)}\Big{)}^{1/2}.$
$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{2,K}^{2}+|\Pi_{h}^{\nabla}u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}\Big{)}^{1/2}|\phi|_{2}.$
$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}h_{K}^{-2}|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{1,K}^{2}\Big{)}^{1/2}|\phi|_{2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{4,K}^{2}(u_{h})\Big{)}^{1/2}|\phi|_{2}.$
@@
For $J_{1}(u_{h},I_{h}\phi)$, by the trace inequality (2.5), the boundedness of $\Pi_{h}^{\nabla}$ and interpolation error estimate (3.8), we have
$J_{1}(u_{h},I_{h}\phi)$ $\lesssim(\sum_{e\in\mathcal{E}_{h}}\eta_{1,e}^{2})^{1/2}\Big{(}\sum_{e\in\mathcal{E}_{h}}|e|^{-1}\Big{\|}\Big{[}\frac{\partial(\Pi_{h}^{\nabla}I_{h}\phi-\phi)}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\Big{)}^{1/2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{1,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$
For $J_{3}(u_{h},I_{h}\phi)$, by the Cauchy-Schwarz inequality, trace inequality (2.5), the inverse inequalities of polynomials and the boundedness of $\Pi_{h}^{\nabla}$, ones give
$J_{3}(u_{h},I_{h}\phi)$ $=\Big{|}\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\nabla}I_{h}\phi}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s\Big{|}\lesssim(\sum_{e\in\mathcal{E}_{h}}\eta_{1,e}^{2})^{1/2}\Big{(}\sum_{K\in\mathcal{T}_{h}}|\Pi_{h}^{\nabla}I_{h}\phi|_{2,K}^{2}\Big{)}^{1/2}$
$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{1,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$ (5.6)